A toroird having a square cross section , 5.20cm on edge , and an inner radius of 16.2cm has 535 turns and carries a current of 813mA calulate the magnetic field inside the toroid at (a) the inner and (b) the outer radius of the toroid

Grade:12th pass

2 Answers

Vikas TU
6 years ago
Using the equation, B = (μoIN/2πr)
The inner radius is r = 16.2 cm, so the field there is
B = (4π 10-7).(0.813).(535)/2π(0.162)
= 5.37 × 10-4 T
 
The outer radius is r = 16.2 + 5.2 = 21.4 cm.
The field there is B = (4π 10-7).(0.813).(535)/2π(0.214)
= 4.06 × 10-4 T
 
Rana Zaman
one year ago
A solenoid that is 95.6 cm long has a radius of 1.9 cm and a winding of 1230 turns; it carries a current of 3.58 A. Calculate the magnitude of the magnetic field inside the solenoid